OpenStudy (anonymous):

Solve by elimination 7x + 3y = 25 -2x - y = -8

OpenStudy (anonymous):

solving by elimination means we need the coefficients on one of the variables to be the same or opposite. so lets multiply the bottom equation -2x - y = -8 by 3 we would get -6x - 3y = -24 now we can literally add the equations together to get rid of the y variable

OpenStudy (anonymous):

7x+3y = 25 + -6x -y =-24 x = 1

OpenStudy (anonymous):

now that we know x = 1, we can plug that back into one of the equations and solve for y so 7(1) + 3y = 25 now just solve for y

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

you're welcome